OpenStudy (anonymous):

Help please Multiply: (2 + square root sign 7) (1+3 square root sign 7)

OpenStudy (anonymous):

\[(2+\sqrt{7})(1+3\sqrt{7})\]

OpenStudy (anonymous):

\[ 23+7\sqrt{7} \]

hero (hero):